OpenStudy (anonymous):

Solving Systems by Substitution y=2x y=-6x+4

OpenStudy (anonymous):

-6x?

OpenStudy (anonymous):

2x=-6+4 2x=-2 x=-2/2 x=-1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case you can just do: \[2x = -6+4\] and solve away!

OpenStudy (anonymous):

Another way to write the system is: 2x = y = -6x + 4 or 2x = -6x + 4. Solve for x, then plug that value into either equation to get y.